OpenStudy (anonymous):

how do you change the expression 4x+1 to be equivalent to 5-4(x-1)?

OpenStudy (anonymous):

4x + 1 5 - 4(x-1) = 5 - 4x +4 = 9-4x 4x + 1 multiply by minus -4x - 1 add 10 9 -4x I am no5 sure if this is what you are lookin for